α \alpha  θ \theta  o o  τ \tau
β \beta  ϑ \vartheta π \pi  υ \upsilon
γ \gamma  γ \gamma  ϖ \varpi  ϕ \phi
δ \delta  κ \kappa  ρ \rho  φ \varphi
ϵ \epsilon  λ \lambda  ϱ \varrho  χ \chi
ε \varepsilon μ \mu  σ \sigma  ψ \psi
ζ \zeta  ν \nu  ς \varsigma ω \omega
η \eta  ξ \xi
Γ \Gamma  Λ \Lambda  Σ \Sigma  Ψ \Psi
Δ \Delta  Ξ \Xi  Υ \Upsilon Ω \Omega
Θ \Theta  Π \Pi  Φ \Phi

Table 1: Greek Letters

± \pm  ∩ \cap  ⋄ \diamond  ⊕ \oplus
∓ \mp  ∪ \cup  △ \bigtriangleup  ⊖ \ominus
× \times ⊎ \uplus  ▽ \bigtriangledown ⊗ \otimes
÷ \div  ⊓ \sqcap  ◁ \triangleleft  ⊘ \oslash
∗ \ast  ⊔ \sqcup  ▷ \triangleright  ⊙ \odot
⋆ \star  ∨ \vee  < \lhd^{b}  ◯ \bigcirc
∘ \circ  ∧ \wedge  > \rhd^{b}  † \dagger
• \bullet \ \setminus ≤ \unlhd^{b}  ‡ \ddagger
· \cdot  ≀ \wr  ≥ \unrhd^{b}  ⨿ \amalg
+ +  − -

b Not predefined in a format based on basefont.tex. Use one of the style options

           oldlfont, newlfont, amsfonts or amssymb.

Table 2: Binary Operation Symbols

≤ \leq  ≥ \geq  ≡ \equiv ⊧ \models
≺ \prec  ≻ \succ  ∼ \sim  ⊥ \perp
⪯ \preceq  ⪰ \succeq  ≃ \simeq | \mid
≪ \ll  ≫ \gg  ≍ \asymp ∥ \parallel
⊂ \subset  ⊃ \supset  ≈ \approx ▷◁ \bowtie
⊆ \subseteq  ⊇ \supseteq    ≅ \cong  ⋈ \Join^{b}
⊏ \sqsubset^{b} ⊐ \sqsupset^{b} ≠ \neq  ⌣ \smile
⊑ \sqsubseteq ⊒ \sqsupseteq  ≐ \doteq ⌢ \frown
∈ \in  ∋ \ni  ∝ \propto = =
⊢ \vdash  ⊣ \dashv  < <  > >
: :

b Not predefined in a format based on basefont.tex. Use one of the style options

           oldlfont, newlfont, amsfonts or amssymb.

Table 3: Relation Symbols

   , ,  ; ;  : \colon . \ldotp · \cdotp

Table 4: Punctuation Symbols

← \leftarrow  ⟵ \longleftarrow  ↑ \uparrow
⇐ \Leftarrow  ⟸ \Longleftarrow  ⇑ \Uparrow
→ \rightarrow  ⟶ \longrightarrow  ↓ \downarrow
⇒ \Rightarrow  ⟹ \Longrightarrow  ⇓ \Downarrow
↔ \leftrightarrow  ⟷ \longleftrightarrow ↕ \updownarrow
⇔ \Leftrightarrow  ⟺ \Longleftrightarrow ⇕ \Updownarrow
↦ \mapsto  ⟼ \longmapsto  ↗ \nearrow
↩ \hookleftarrow  ↪ \hookrightarrow  ↘ \searrow
↼ \leftharpoonup  ⇀  \rightharpoonup  ↙ \swarrow
↽ \leftharpoondown  ⇁  \rightharpoondown  ↖ \nwarrow
⇀ ↽ \rightleftharpoons ⇝  \leadsto^{b}

b Not predefined in a format based on basefont.tex. Use one of the style options

                oldlfont, newlfont, amsfonts or amssymb.

Table 5: Arrow Symbols

   . . . \ldots · · · \cdots     ⋮ \vdots  ... \ddots
   ℵ \aleph ′  \prime  ∀ \forall  ∞ \infty
   h̄ \hbar ∅ \emptyset ∃ \exists  □ \Box^{b}
   ı  \imath ∇ \nabla  ¬ \neg  ◇ \Diamond^{b}
   ȷ \jmath √ \surd  ♭ \flat  △ \triangle
   ℓ \ell  ⊤ \top  ♮ \natural  ♣ \clubsuit
   ℘ \wp  ⊥ \bot  ♯ \sharp  ♢ \diamondsuit
   ℜ \Re  ∥ \|  \ \backslash ♡ \heartsuit
   ℑ \Im     ^{}\angle  ∂ \partial  ♠ \spadesuit
   ℧ \mho^{b} .  .  | |

b Not predefined in a format based on basefont.tex. Use one of the style options

                oldlfont, newlfont, amsfonts or amssymb.

Table 6: Miscellaneous Symbols

            ∑ \sum  ⋂ \bigcap  ⨀ \bigodot
            ∏ \prod  ⋃ \bigcup  ⨂ \bigotimes
            ∐ \coprod ⨆ \bigsqcup ⨁ \bigoplus
            ∫ \int  ⋁ \bigvee  ⨄ \biguplus
            ∮ \oint  ⋀ \bigwedge

Table 7: Variable-sized Symbols

    \arccos  \cos   \csc  \exp  \ker     \limsup  \min  \sinh
    \arcsin  \cosh  \deg  \gcd  \lg      \ln      \Pr   \sup
    \arctan  \cot   \det  \hom  \lim     \log     \sec  \tan
    \arg     \coth  \dim  \inf  \liminf  \max     \sin  \tanh

Table 8: Log-like Symbols

     ( (  ) )  ↑ \uparrow  ⇑ \Uparrow
     [ [  ] ]  ↓ \downarrow  ⇓ \Downarrow
     { \{  } \}  ↕ \updownarrow ⇕ \Updownarrow
     ⌊ \lfloor ⌋ \rfloor ⌈ \lceil  ⌉ \rceil
     ⟨ \langle ⟩ \rangle / /  \ \backslash
     | |  ∥ \|

Table 9: Delimiters

} \rmoustache  { \lmoustache  } \rgroup     { \lgroup
| \arrowvert   ‖ \Arrowvert   | \bracevert

Table 10: Large Delimiters

â \hat{a}    á \acute{a}  ā \bar{a}  ȧ \dot{a}   ă \breve{a}
ǎ \check{a}  à \grave{a}  a⃗ \vec{a}  ä \ddot{a}  ã \tilde{a}

Table 11: Math mode accents

ãbc  \widetilde{abc}      âbc  \widehat{abc}
\overleftarrow{abc}  \overleftarrow\{abc\}  \overrightarrow{abc}  \overrightarrow\{abc\}
\overline{abc}  \overline\{abc\}       \underline{abc}  \underline\{abc\}
\overbrace{abc} \overbrace\{abc\}      \underbrace{abc} \underbrace\{abc\}
\sqrt{abc} \sqrt\{abc\}            \sqrt[n]{abc} \sqrt[n]\{abc\}
f′  f'                    \frac{abc}{xyz}  \frac\{abc\}\{xyz\}

Table 12: Some other constructions